# Development of large-volume 130TeO2 bolometers for the CROSS $2\beta$ decay search experiment

### v1

@@ Abstract
We report on the development of thermal detectors based on large-size tellurium dioxide crystals (45$\times$45$\times$45 mm), containing tellurium enriched in 130Te to about 91%, for the CROSS double-beta decay experiment. A powder used for the crystals growth was additionally purified by the directional solidification method, resulting in the reduction of the concentration of impurities by a factor 10, to a few ppm of the total concentration of residual elements (the main impurity is Fe). The purest part of the ingot (the first $\sim$200 mm, about 80% of the total length of the cylindrical part of the ingot) was determined by scanning segregation profiles of impurities (both on the surface and in center of the ingot) and used for the 130TeO2 powder production with no evidence of re-contamination. The crystal growth was verified with precursors produced from powder with natural Te isotopic composition, and two small-size (20$\times$20$\times$10 mm) samples were tested at a sea-level laboratory showing high bolometric and spectrometric performance together with acceptable 210Po content (below 10 mBq/kg). This growth method was then applied for the production of six large cubic 130TeO2 crystals for the CROSS experiment and 4 of them were taken randomly to be characterized at the Canfranc underground laboratory, in the CROSS-dedicated low-background cryogenic facility. Two 130TeO2 samples were coated with a thin metal film in form of Al layer (on 4 sides) or AlPd grid (on a single side) to investigate the possibility to tag surface events by pulse-shape discrimination. Similarly to the small natural precursors, large-volume 130TeO2 bolometers show high performance and even better internal purity (210Po activity $\sim$ 1 mBq/kg, while activities of 228Th and 226Ra are below 0.01 mBq/kg), satisfying requirements for the CROSS and, potentially, next-generation experiments.
@@ 1 Introduction
Among the candidates studied, 130Te stands out as one of the few extensively researched $2\beta$-active isotopes, thanks to: a) a relatively high transition energy ($Q_{2\beta}$ = 2527 keV [7]); b) a considerable presence in natural tellurium ($\delta$ = 34% [8]); c) the possibility to grow large-volume, high-quality, and radiopure tellurium dioxide (TeO2) crystals, i.e. a dielectric material with a high tellurium content suitable for the use in low-temperature detectors of rare events [9, 10, 11, 12]. Thanks to the above-listed features, TeO2 is among a few materials with the longest history of applications to $2\beta$ searches. Indeed, initial attempts were made in the early 1990s using small samples. Subsequently, these efforts were expanded to larger arrays in experiments such as MiBETA, Cuoricino, and CUORE-0. Currently, this research has reached its pinnacle with the CUORE ton-scale experiment [13], which is operating at the Gran Sasso underground laboratory (LNGS) in Italy. (The evolution of TeO2-based bolometric $2\beta$ search experiments is overviewed in [14].)
@@
All past and present TeO2-based $2\beta$ experiments exploited and currently used crystals with natural Te content, with the exception of two samples enriched in 130Te used in addition to natural ones in the MiBETA [15] and Cuoricino [16] experiments. Due to poor performance and radiopurity of enriched samples compared to natural crystals, as well as taking into account the still high natural isotopic abundance of 130Te, 130Te-enriched crystals were abandoned from the CUORE program. However, the interest in 130TeO2 has raised again in view of the CUORE upgrade [17], particularly in the CUPID program [18, 19]. Two new enriched samples have been produced and tested at LNGS showing high performance and low radioactivity [20]. Due to challenging particle identification with TeO2 bolometers [21], mainly exploiting the detection of $\gamma(\beta)$-induced Cherenkov radiation proposed in [22], and difficulties in further reduction of $\gamma$ background in the CUORE cryostat, CUPID has selected lithium molybdate (Li2MoO4) scintillating crystals [23]. This compound allows for a more efficient scintillation-assisted particle identification and a lower $\gamma$ background, by moving the region of interest ($Q_{2\beta}$ of 100Mo is 3034 keV [7]) above the end-point of natural radioactivity (2615 keV $\gamma$ of 208Tl from the 232Th family).
@@
Nevertheless, 130Te remains one of a few $2\beta$-active isotopes highly suitable for multi-isotope bolometric study [24]. Furthermore, TeO2 material has a few advantages compared with Li2MoO4: a) higher detection efficiency for the same volume (higher density and effective atomic number); b) no hygroscopic properties; c) two orders of magnitude longer half-life of the ordinary $2\nu 2\beta$ decay, which makes the background contribution from random coincidences of the associated events negligible for 130Te-enriched thermal detectors, whereas this is the major component of background expected in CUPID with 100Mo-enriched scintillating bolometers [25]. Taking into account the possibility to improve the present, already extremely low, $\gamma$ background in CUORE by even more careful selection of external screens with lower contamination and/or by applying innovative techniques for active suppression of backgrounds induced by $\gamma$ (e.g. proposed by BINGO [26]) and $\alpha$ (e.g. by detection of $\gamma$($\beta$)-induced Cherenkov radiation from TeO2 [22]), the choice of 130TeO2 for future high-sensitivity $2\beta$ searches is still attractive and viable.
@@
With this in mind, TeO2 together with Li2MoO4 detector materials have been chosen for the implementation of the CROSS program on the development of metal-coated bolometers with the capability of pulse-shape discrimination between bulk and near surface interactions [27, 28]. The validation of the CROSS technology for active rejection of surface-originated $\alpha$ and $\beta$ backgrounds is planned to be realized in a small-scale demonstrator with tens of Li2100MoO4 and a few 130TeO2 bolometers. For the sake of completeness, we refer to early attempts in the development of active background rejection for bolometric $2\beta$ detectors, including TeO2, which are summarized in [21] and we underline a couple of currently ongoing projects on this topic. As in CROSS, TeO2 and Li2MoO4 are both included in the BINGO project, which is developing an innovative detector assembly system with high-performance bolometric photodetectors and a cryogenic active veto based on BGO crystals with bolometric photodetection [26]. Recently, the BINGO detector system with the nylon wire serving to hold a similar size Li2MoO4 crystal (twice lighter than TeO2) have been tested underground in the CROSS cryostat and good detector performance has been reported [29]. Also, the idea of scintillating polymer-based holder allowing active background rejection of radioactivity induced by close materials of the detector structure has been proposed. First prototypes with small and large TeO2 bolometers were tested in such a 3D-printed detector holder [30, 31], as a part of R&D for next-generation $2\beta$ experiments with this detector material.
@@ 2.1 Initial 130Te-enriched material
The initial tellurium powder enriched in 130Te isotope to around 93% (provided by the University of South Carolina) was assayed as received from the customer for impurities and oxygen content. Using the inductively coupled plasma mass-spectrometry (ICP-MS), the total concentration of impurities excluding oxygen were measured at 25 ppm; with the duplicate sample, a consistent result of 27 ppm was obtained. The detected impurities are listed in table 1. The oxygen content is found on the level of 2%.
Table 1: Impurities detected in samples which were produced during an R&D on purification of the 130Te-enriched material. We quote results for initial, as-purchased, raw material, after purification (average for pure section of the ingot produced by solidification, see text), and after the conversion to a 130TeO2 powder. No results are given for cases when the signal is below the detection limit. Uncertainties are at 20% level.
@@ 2.2 Purification by the solidification method
The purification of the enriched material has been performed in Canada (5N company, Garand, Montreal, Quebec). At the initial step — hydrogenation — a deoxidized ingot was produced from the tellurium oxide powder. The material was remelted under a stream of hydrogen to obtain a solid ingot, as illustrated in figure 1.
@@
Then the directional solidification method was applied. The tellurium ingot from the hydrogenation step was remelted under a partial atmosphere of hydrogen and solidified directionally to get impurity segregation and to obtain a material with a 5N purity. After the process, the ingot was cut at 6 positions over the length, as shown in figure 2, and each position was sampled in the center and on the surface of the ingot. Each sample from the different positions in the ingot was assayed to get segregation profiles, showing signals of impurities above the detection limit at positions of 180–210 mm. This is illustrated in figure 3 for the main impurity (Fe) of the 130Te-enriched material. Therefore, the impure section above 180 mm was removed, while the rest of the ingot was ground and the powder was homogenized for a final sampling. The final assay of the purified tellurium, done using ICP-MS measurements, is presented in table 1. The total impurity in the 130Te-enriched powder is detected at 3 ppm, showing no hint of a re-contamination of the material to be used for the crystal growth.
Figure 3: Segregation profile of iron in samples taken from the center (blue) and on the surface (orange) of the 130Te-enriched tellurium ingot produced by the directional solidification.
@@ 2.4 Crystal growth
We identified a company in the US, Gooch & Housego (Clevelend, Ohio), for the Czochralski growth of large-volume (about 90 cm3) 130Te-enriched TeO2 crystals to be included in the CROSS demonstrator. In order to validate the crystal growth process, two natural TeO2 samples (24 g mass each) were produced by this company from non-purified materials and tested as Cherenkov-radiation-emitting bolometers in Orsay to demonstrate acceptable bolometric performance and the level of internal 210Po contamination (less than 10 mBq/kg), as detailed in the next section. After the validation of the TeO2 crystal producer, we proceeded with the growth of 130Te-enriched crystals and six 130TeO2 samples 45$\times$45$\times$45 mm each were fabricated using purified starting materials.
@@ 3.1 Above-ground test of small-size natural precursors
On the top of the TeO2 assembly we placed a low-threshold optical bolometer [35] to attempt to detect the tiny Cherenkov radiation induced by $\beta$ / $\gamma$ / $\mu$ interactions in the TeO2 bolometers [36], thus allowing for selection and study of $\alpha$ events. This photodetector (NTLLD2 in [35]) was made of an electronic grade pure Ge slab ($\oslash 44\times 0.18$ mm) and instrumented with a smaller size NTD-Ge sensor epoxy-glued on the wafer surface. The Ge wafer was coated by a 70-nm SiO layer to reduce the reflectivity of the surface and thus to improve photon collection. A set of aluminum concentric electrodes were deposited on the Ge surface and interconnected with Al wires, allowing to apply a voltage bias and to amplify thermal signals exploiting the Neganov-Trofimov-Luke effect in semiconductors. The Ge slab was kept in a Cu housing with the help of sapphire balls ($\varnothing 1.5$ mm) and polypropylene supporting elements; the NTD-Ge sensor was polarized through Au bonding wires. The mounted optical bolometer is depicted in figure 5 (Right). The photodetector was facing the TeO2 crystals by the side carrying the Al electrodes, while the opposite side was facing another bolometer (ZnO). Its recently published characterization [34] contains more details on the experiment and results of the light detector performance, which are only briefly described below.
@@
Measurements were carried out with a 232Th calibration source for 24 h and without the use of additional radioactive sources, the nominally background data were taken for a total of 166 hours. Due to a comparatively high environmental $\gamma$-radioactivity (mainly from decays of the 226Ra sub-chain [41]), and thanks to high spectroscopic properties of the TeO2 bolometers, background data can be easily calibrated because of the presence of many sharp peaks in a wide energy range as seen in figure 6. We observed that both TeO2 thermal detectors show excellent energy resolution, comparable to results achieved with bare (i.e. non-coated) samples of the same-size used in the CROSS program [27, 42]. The detectors are characterized by high sensitivity defined as the voltage signal per unit of deposited energy (0.2–0.5 $\mu$V/keV) and a low noise represented by the baseline width at the optimum filter output (0.25–0.60 eV RMS). As for the time properties of the thermal response, both TeO2 bolometers show typically slow signals with the ascending part of about 12 ms and with the descending part about 60 ms. The performances of the TeO2 thermal detectors are summarized in table 3.
Table 3: Performance of two 24-g TeO2 bolometers and a Ge light detector exploiting the signal amplification based on the Neganov-Trofimov-Luke effect (60 V bias on Al electrode). We report the working point of the NTD thermistor (resistance $R_{NTD}$ at a given current $I_{NTD}$), pulse shape time constants (rise- and decay-time parameters, $\tau_{r}$ and $\tau_{d}$), detector sensitivity represented by the signal voltage amplitude per unit of the deposited energy ($A_{s}$), RMS baseline width (RMSnoise).
@@
Figure 7: Light signals detected by the Ge optical bolometer versus heat signals measured by the 24-g TeO2 bolometer (TeO-81) over 96 h of exposure to natural radioactivity at a sea-level laboratory. The major part of events below $\sim$2.6 MeV corresponds to $\gamma$($\beta$) interactions in the crystal medium, while muon-induced events represent dominant contribution above this energy. A structure of high-energy TeO2-detected events which exhibit no light signal correspond to $\alpha$ decays of 210Po inside the TeO2 detector medium, indicating an internal contamination by this radionuclide.
During the TeO2 detector calibrations we kept the Al electrodes of the Ge bolometer unpolarized, thus allowing the detection of Te X-rays induced by environmental radioactivity in the crystals, which was used to calibrate the maximum distribution of the muon-induced events [34]. Knowing the energy scale of the Ge bolometer, we found a factor $\sim$10 improvement of the signal-to-noise ratio by applying a 60 V bias on the Al electrode, resulting to a high performance reported in table 3 and in [34]. Thanks to the low noise achieved with the Ge bolometer, we were able to detect the TeO2-emitted Cherenkov light, providing highly efficient particle identification as illustrated in figure 7. As one can see in figure 7, most of the events populate a band including $\beta$-, $\gamma$- and $\mu$-induced interactions, which are associated with a registered Cherenkov light of the order of 40 eV per MeV energy deposited in the TeO2 bolometer. The population of $\alpha$ events, which do not produce the Cherenkov radiation (due to energy threshold of 400 MeV), is also observed in the data (see figure 7). Moreover, these $\alpha$-events have even negative values of light signals due to a small cross-talk, also seen in the data for heater-induced events (not shown in figure 7). The observed $\alpha$ events are mainly ascribed to a 210Po contamination in the crystal bulk corresponding to an activity of 2.8(4) mBq/kg and 9.1(8) mBq/kg in the samples TeO-63 and TeO-81 respectively. This contamination is typical for TeO2, as well as for different crystal scintillators [43], but does not originate by the 210Pb content as for many other materials. It is worth noting that we detected only tens of 210Po events, while the presence of other $\alpha$-active radionuclides from U/Th chains is not evident, and thus their activities are expected to be at least an order of magnitude lower. Therefore, one can conclude that the radiopurity of both samples is very high, especially keeping in mind that the starting material did not pass an extra purification.

### v2

@@ Abstract
We report on the development of thermal detectors based on large-size tellurium dioxide crystals (45 $\times$ 45 $\times$ 45 mm), containing tellurium enriched in 130Te to about 91%, for the CROSS double-beta decay experiment. A powder used for the crystals growth was additionally purified by the directional solidification method, resulting in the reduction of the concentration of impurities by a factor 10, to a few ppm of the total concentration of residual elements (the main impurity is Fe). The purest part of the ingot (the first $\sim$200 mm, about 80% of the total length of the cylindrical part of the ingot) was determined by scanning segregation profiles of impurities and used for the 130TeO2 powder production with no evidence of re-contamination. The crystal growth was verified with precursors produced from powder with natural Te isotopic composition, and two small-size (20 $\times$ 20 $\times$ 10 mm) samples were tested at a sea-level laboratory showing high bolometric and spectrometric performance together with acceptable 210Po content (below 10 mBq/kg). This growth method was then applied for the production of six large cubic 130TeO2 crystals and 4 of them were taken randomly to be characterized at the Canfranc underground laboratory, in the CROSS-dedicated low-background cryogenic facility. Two 130TeO2 samples were coated with a thin, $O$(100 nm), metal film in form of Al layer (on 4 sides) or AlPd grid (on a single side) to investigate the possibility to tag surface events by pulse-shape discrimination. Similarly to the small natural precursors, large-volume 130TeO2 bolometers show high performance and even better internal purity (210Po activity $\sim$ 1 mBq/kg, while activities of 228Th and 226Ra are below 0.01 mBq/kg), satisfying requirements for the CROSS and, potentially, next-generation experiments.
@@ 1 Introduction
Among the candidates studied, 130Te stands out as one of the few extensively researched $2\beta$-active isotopes, thanks to: i) a relatively high transition energy ($Q_{2\beta}$ = 2527 keV [7]); ii) a considerable presence in natural tellurium ($\delta$ = 34% [8]); iii) the possibility to grow large-volume, high-quality, and radiopure tellurium dioxide (TeO2) crystals, i.e. a dielectric material with a high tellurium content suitable for the use in low-temperature detectors of rare events [9, 10, 11, 12]. Thanks to the above-listed features, TeO2 is among a few materials with the longest history of applications to $2\beta$ searches. Indeed, initial attempts were made in the early 1990s using small samples. Subsequently, these efforts were expanded to larger arrays in experiments such as MiDBD, Cuoricino, and CUORE-0. Currently, this research has reached its pinnacle with the CUORE ton-scale experiment [13], which is operating at the Gran Sasso underground laboratory (LNGS) in Italy. (The evolution of TeO2-based bolometric $2\beta$ search experiments is overviewed in [14].)
@@
All past and present TeO2-based $2\beta$ experiments exploited and currently used crystals with natural Te content, with the exception of two samples enriched in 130Te used in addition to natural ones in the MiDBD [15] and Cuoricino [16] experiments. Due to poor performance and radiopurity of enriched samples compared to natural crystals, as well as taking into account the still high natural isotopic abundance of 130Te, 130Te-enriched crystals were abandoned from the CUORE program. However, the interest in 130TeO2 has raised again in view of the CUORE upgrade [17], particularly in the CUPID program [18, 19]. Two new enriched samples have been produced and tested at LNGS showing high performance and low radioactivity [20]. Due to challenging particle identification with TeO2 bolometers [21], mainly exploiting the detection of $\gamma(\beta)$-induced Cherenkov radiation proposed in [22], and difficulties in further reduction of $\gamma$ background in the CUORE cryostat, CUPID has selected lithium molybdate (Li2MoO4) scintillating crystals [23]. This compound allows for a more efficient scintillation-assisted particle identification and a lower $\gamma$ background, by moving the region of interest ($Q_{2\beta}$ of 100Mo is 3034 keV [7]) above the end-point of the most intense natural $\gamma$ radioactivity (2615 keV $\gamma$ of 208Tl from the 232Th family).
@@
Nevertheless, 130Te remains one of a few $2\beta$-active isotopes highly suitable for multi-isotope bolometric study [24]. Furthermore, TeO2 material has a few advantages compared with Li2MoO4: i) higher detection efficiency for the same volume (higher density and effective atomic number); ii) no hygroscopic properties; iii) two orders of magnitude longer half-life of the ordinary $2\nu 2\beta$ decay, which makes the background contribution from random coincidences of the associated events negligible for 130Te-enriched thermal detectors, whereas this is the major component of background expected in CUPID with 100Mo-enriched scintillating bolometers [25]. Taking into account the possibility to improve the present, already extremely low, $\gamma$ background in CUORE by even more careful selection of external screens with lower contamination and/or by applying innovative techniques for active suppression of backgrounds induced by $\gamma$ (e.g. proposed by BINGO [26]) and $\alpha$ (e.g. by detection of $\gamma$($\beta$)-induced Cherenkov radiation from TeO2 [22]), the choice of 130TeO2 for future high-sensitivity $2\beta$ searches is still attractive and viable.
@@
With this in mind, TeO2 together with Li2MoO4 detector materials have been chosen for the implementation of the CROSS program on the development of metal-coated bolometers with the capability of pulse-shape discrimination between bulk and near surface interactions [27, 28]. The validation of the CROSS technology for active rejection of surface-originated $\alpha$ and $\beta$ backgrounds is planned to be realized in a small-scale demonstrator with tens of Li2100MoO4 and a few 130TeO2 bolometers. For the sake of completeness, we refer to early attempts in the development of active background rejection for bolometric $2\beta$ detectors, including TeO2, which are summarized in [21] and we underline a couple of currently ongoing projects on this topic. As in CROSS, TeO2 and Li2MoO4 are both included in the BINGO project, which is developing an innovative detector assembly system with high-performance bolometric photodetectors and a cryogenic active veto based on bismuth germanate crystals with bolometric photodetection [26]. Recently, the BINGO detector system with the polyamide wire serving to hold a similar size Li2MoO4 crystal (twice lighter than TeO2) have been tested underground in the CROSS cryostat and good detector performance has been reported [29]. Also, the idea of scintillating polymer-based holder allowing active background rejection of radioactivity induced by close materials of the detector structure has been proposed [30]. First prototypes with small and large TeO2 bolometers were tested in such a 3D-printed detector holder [31, 30], as a part of R&D for next-generation $2\beta$ experiments with this detector material.
@@ 2.1 Initial 130Te-enriched material
The initial tellurium powder enriched in 130Te isotope to around 93% (provided by the University of South Carolina) was assayed as received from the customer for impurities and oxygen content. Using the inductively coupled plasma mass-spectrometry (ICP-MS), the total concentration of impurities excluding oxygen was measured at 25 ppm; with the duplicate sample, a consistent result of 27 ppm was obtained. The detected impurities are listed in table 1. The oxygen content is found on the level of 2%.
Table 1: ICP-MS-detected impurities in samples which were produced during an R&D on purification of the 130Te-enriched material. We quote results for initial (as-purchased) raw material, after purification (average for pure section of the ingot produced by solidification, see text), and after the conversion to a 130TeO2 powder. Uncertainties are at 20% level. No results are given for cases when the signal is below the detection limit (including the following elements: Ag, As, B, Bi, Cd, Co, Mo, P, Sb, Se, and V).
@@ 2.2 Purification by the solidification method
The purification of the enriched material has been performed in Canada (5N company, Garand, Montreal, Quebec). At the initial step — hydrogenation process [32] — a deoxidized ingot was produced from the tellurium oxide powder. The material was remelted under a stream of hydrogen to obtain a solid ingot, as illustrated in figure 1.
@@
Then the directional solidification method [33] was applied. The tellurium ingot from the hydrogenation step was remelted under a partial atmosphere of hydrogen and solidified directionally to get impurity segregation and to obtain a material with a 5N purity. After the process, the ingot was cut at 6 positions over the length, as shown in figure 2, and each position was sampled in the center and on the surface of the ingot. Each sample from the different positions in the ingot was assayed to get segregation profiles, showing signals of impurities above the detection limit at positions of 180–210 mm. This is illustrated in figure 3 for the main impurity (Fe) of the 130Te-enriched material; the segregation profiles of other impurities show the same trend. Therefore, the impure section above 180 mm was removed, while the rest of the ingot was ground and the powder was homogenized for a final sampling. The final assay of the purified tellurium, done using ICP-MS measurements, is presented in table 1. The total impurity in the 130Te-enriched powder is detected at 3 ppm, showing no hint of a re-contamination of the material to be used for the crystal growth.
Figure 3: Segregation profile of iron (the main impurity) in samples taken from the center (blue) and on the surface (orange) of the 130Te-enriched tellurium ingot produced by the directional solidification.
@@ 2.4 Crystal growth
We identified a company in the US, Gooch & Housego (Clevelend, Ohio), for the Czochralski growth of large-volume (about 90 cm3) 130Te-enriched TeO2 crystals to be included in the CROSS demonstrator. In order to validate the crystal growth process, two natural TeO2 samples (4 cm3 each) were produced by this company from non-purified materials and tested as Cherenkov-radiation-emitting bolometers in Orsay to demonstrate acceptable bolometric performance and the level of internal 210Pb contamination (less than 10 mBq/kg), as detailed in the next section. After the validation of the TeO2 crystal producer, we proceeded with the growth of 130Te-enriched crystals and six 130TeO2 samples 45 $\times$ 45 $\times$ 45 mm each were fabricated using purified starting materials.
@@ 3.1 Above-ground test of small-size natural precursors
On the top of the TeO2 assembly we placed a low-threshold optical bolometer [37] to attempt to detect the tiny Cherenkov radiation induced by $\beta$ / $\gamma$ / $\mu$ interactions in the TeO2 bolometers [38], thus allowing for selection and study of $\alpha$ events. This photodetector (NTLLD2 in [37]) was made of an electronic grade pure Ge slab ($\oslash 44\times 0.18$ mm) and instrumented with a smaller size NTD-Ge sensor epoxy-glued on the wafer surface. The Ge wafer was coated by a 70-nm SiO layer to reduce the reflectivity of the surface and thus to improve photon collection. A set of aluminum concentric electrodes were deposited on the Ge surface and interconnected with Al wires, allowing to apply a voltage bias and to amplify thermal signals exploiting the Neganov-Trofimov-Luke effect in semiconductors [39, 40]. The Ge slab was kept in a Cu housing with the help of sapphire balls ($\varnothing 1.5$ mm) and polypropylene supporting elements; the NTD-Ge sensor was polarized through Au bonding wires. The mounted optical bolometer is depicted in figure 5 (Right). The photodetector was facing the TeO2 crystals by the side carrying the Al electrodes (to get more efficient amplification of photon-induced thermal signals), while the opposite side was facing another bolometer (ZnO). Its recently published characterization [36] contains more details on the experiment and results of the light detector performance, which are only briefly described below.
@@
Measurements were carried out with a 232Th calibration source for 24 h and without the use of additional radioactive sources; the nominally background data were taken for a total of 166 hours. Due to a comparatively high environmental $\gamma$-radioactivity (mainly from decays of the 226Ra sub-chain [45]), and thanks to high spectroscopic properties of the TeO2 bolometers, background data can be easily calibrated because of the presence of many sharp peaks in a wide energy range as seen in figure 6. We observed that both TeO2 thermal detectors show excellent energy resolution, comparable to results achieved with bare (i.e. non-coated) samples of the same-size used in the CROSS program [27, 46]. The detectors are characterized by high sensitivity defined as the voltage signal per unit of deposited energy (0.2–0.5 $\mu$V/keV) and a low noise represented by the baseline width at the optimum filter output (0.25–0.60 eV RMS). As for the time properties of the thermal response, both TeO2 bolometers show typically slow signals with the ascending part of about 12 ms and with the descending part about 60 ms. The performances of the TeO2 thermal detectors are summarized in table 3.
Table 3: Performance of two 24-g TeO2 bolometers and a Ge light detector exploiting the signal amplification based on the Neganov-Trofimov-Luke effect (60 V bias on Al electrode). We report the working point of the NTD thermistor (resistance $R_{NTD}$ at a given current $I_{NTD}$), pulse-shape time constants (rise- and decay-time parameters, $\tau_{r}$ and $\tau_{d}$), detector sensitivity represented by the signal voltage amplitude per unit of the deposited energy ($A_{s}$; estimated with 609-keV $\gamma$ quanta), and RMS baseline width (RMSnoise).
@@
Figure 7: Light signals detected by the Ge optical bolometer versus heat signals measured by the 24-g TeO2 bolometer (TeO-81) over 96 h of exposure to natural radioactivity at a sea-level laboratory. The major part of events below $\sim$2.6 MeV corresponds to $\gamma$($\beta$) interactions in the crystal medium, while muon-induced events represent dominant contribution above this energy. A structure of high-energy TeO2-detected events which exhibit no light signal correspond to $\alpha$ decays of 210Po inside the TeO2 detector medium, indicating an internal contamination by 210Po(210Pb). The energy scale of the detector, calibrated using 609-keV $\gamma$ quanta of 214Bi, exhibits an evident non-linearity, which results to a few % miscalibration of the 5407-keV $\alpha$ peak of 210Po (see discussion in Sec. 3.2).
During the TeO2 detector calibrations we kept the Al electrodes of the Ge bolometer unpolarized, thus allowing the detection of Te X-rays induced by environmental radioactivity in the crystals, which was used to calibrate the maximum distribution of the muon-induced events [36]. Knowing the energy scale of the Ge bolometer, we found a factor $\sim$10 improvement of the signal-to-noise ratio by applying a 60 V bias on the Al electrode, resulting to a high performance reported in table 3 and in [36].
Thanks to the low noise achieved with the Ge bolometer, we were able to detect the TeO2-emitted Cherenkov light, providing highly efficient particle identification as illustrated in figure 7. As one can see in figure 7, most of the events populate a band including $\beta$-, $\gamma$- and $\mu$-induced interactions, which are associated with a registered Cherenkov light of the order of 40 eV per MeV energy deposited in the TeO2 bolometer. The population of $\alpha$ events, which do not produce the Cherenkov radiation (due to kinematic energy threshold of 400 MeV), is also observed in the data (see figure 7). Moreover, these $\alpha$-events have even negative values of light signals due to a small cross-talk, also seen in the data for heater-induced events (not shown in figure 7).
The observed $\alpha$ events are mainly ascribed to decays of 210Po ($Q_{\alpha}$ = 5407.5 keV, $T_{1/2}$ = 138.376 d) from 210Po/210Pb contaminations in the crystal bulk corresponding to an activity of 2.8(4) mBq/kg and 9.1(8) mBq/kg in the samples TeO-63 and TeO-81 respectively. The 210Po $\alpha$ activity is typical for different scintillators [47] and mostly related to the 210Pb content in materials. In case of TeO2 crystals, the chemical affinity between Te and Po could result to a less efficient segregation of Po isotopes, including 210Po, during the TeO2 powder purification and crystal growth. It is worth noting that we detected only tens of 210Po events, while the presence of other $\alpha$-active radionuclides from U/Th chains is not evident, and thus their activities are expected to be at least an order of magnitude lower. Therefore, one can conclude that the radiopurity of both samples is very high, especially keeping in mind that the starting material did not pass an extra purification.